# Exploring the interaction via femtoscopic study

### v1

@@ Abstract
Very recently the Wood-Saxon (WS) type interaction in the single-folding potential approach are constructed to simulate the potentials. One set of the potentials are based on the first principle HAL QCD interactions in channel, and in another set, the -meson–nucleus potentials were calculated by employing the quark-meson coupling (QMC) model. By utilizing these two set of potentials, the two-particle momentum correlation of in high-energy heavy ion collisions is explored. The numerical results show that the correlation functions at small source size (high density nuclear medium) depends on the employed potential model. Also, the correlation functions are obtained within the Lednicky-Lyuboshits (LL) formalism. For small source size, it is found that, the LL formula returns significantly different values due to the large interaction range of the potential.
@@ I Introduction
Although considerable experimental and theoretical [1, 2, 3, 4, 5, 6] efforts have been done in the last few decades, the possible existence of -mesic bound states with nucleons (N) is still a challenging field of research. The essential basis for the studying of such a system is an exact information about the strong interaction.
@@
Up to now, no evidence of a bound state has been found, experimentally, since the standard invariant mass measurements of its decay products are disputing and the measurement of the scattering parameters are restricted to spin averaged quantities. A list of measurements that have been made for absorption off different nuclear targets is given in [7], and the authors concluded that the landscape of experimental results for the interaction is not convincingly in agreement and new information are required. One method for studying the hadron-hadron interaction that is hard to investigate in scattering experiment is measuring the momentum correlation functions in high-energy collisions [8]. It can provides information on both the effective emission source and the interaction potential. Very recently, the ALICE collaboration measured the correlation function of proton- in heavy-ion collisions [9], together by indicating a bound state using two-particle correlation functions [7] request the proton bound state hypothesis.
@@
To go one step further, the femtoscopic analysis of hadron-nuclei correlation functions could play a crucial role in understanding the structure and dynamics of the atomic nuclei [10, 11]. Accordingly, the correlation functions are investigated for some hadron-deuteron systems like [12, 10, 13], [14, 13], [15], [16], and [17, 18]. Furthermore, the momentum correlation between [19], [20] and [21] have recently been studied theoretically. From experimental point of view, the momentum correlation between (anti)protons and (anti)deuterons measured in pp collisions at TeV with ALICE is studied for the first time [22] and the first measurement of correlation with GeV Au+Au collisions is done by STAR Collaborations [23]. In addition, there is a research plan at the Thomas Jefferson National Accelerator Facility (JLab), to investigate the binding of (and ) to [4].
@@
A recent HAL QCD simulations of the potential in channel (by the notation where and are the total spin, orbital angular momentum and total angular momentum, respectively) are performed [24]. The simulation is done on the -flavor with nearly physical quark masses MeV on a large lattice volume of at the imaginary-time slices where fm is the lattice spacing. In Ref. [25] the potential was obtained through a single folding of the HAL QCD interaction in the channel with the nuclear matter distribution of . Moreover, a Wood-Saxon (WS) type interaction was built to simulate the potential, given in [4], based on an effective Lagrangian approach where the nuclear density distributions, as well as the in-medium and meson masses, were consistently calculated in the frame work of quark-meson coupling (QMC) model [26]. In Ref. [25], the authors concluded that there were qualitatively a good agreement between binding energies for both types of the interactions.
@@ II Features of potentials
In the single folding approach the effective nuclear potential can be approximated as
where is potential between the particle at and the nucleon at r [27, 28, 29]. And is the nucleon density distribution in -particle at a distance from its center-of-mass. The detailed calculations of how to obtain the is given by Filikhin et al in [25]. Here, I just mention the main part. As follows from Ref. [30], Filikhin et al. chose the simple Gaussian matter density distribution model as
@@
In Ref. [25], for phenomenological application and calculation of observables, such as scattering phase shifts and binding energies, the is fitted to a Wood-Saxon form using the function that given by Eq. (3) (motivated by common Dover-Gal model of potential [31]) with three parameters, and ,
represents strength of the interaction and known as the depth parameter, where is the mass number of the nucleus and defines the surface diffuseness. The obtained value of the parameters of the Eq. (3) correspond to these three rms radius are given in the table 1 and illustrated in Fig. 1. Note that concrete parameterizations are taken straight from Ref. [25].
@@
In Ref. [4], the -meson–nuclear potentials were calculated employing a local density approximation, by evaluating the loop contribution in the self-energy. The nuclear density distributions, as well as the in-medium and meson masses, were consistently calculated in the frame work of quark-meson coupling (QMC) model [26]. Furthermore, they investigated the sensitivity of the results to the cutoff parameter in the form factor at the vertex appearing in the -meson self-energy. The resultant potential is given for three values of the cutoff parameter, i.e., and GeV.
Accordingly, Filikhin et al. in Ref. [25], constructed a Wood-Saxon (WS) type potentials to simulate potentials for and GeV. In Fig. 1, for comparison, I just represent the potentials, denoted as by using the parameters given in Ref. [25] directly, for three values of the cutoff parameter and GeV. In table 1, the fit parameters of potential in Eq. (3) based on HAL QCD (at lattice Euclidean time ) and quark-meson coupling model of potentials, are presented.
@@
By employing potentials as input, the Schrödinger equation were solved in the infinite volume to extract binding energy and scattering observables from the asymptotic behavior of the wave function. Fig. 2 shows normalized phase shifts calculated as functions of the magnitude of the relative momentum ( is the reduced mass of system) with model of HAL QCD for the three values of rms and QMC potentials for the three values of . The phase shift behaviour for the all case represents an attractive interaction even to form a bound state. The scattering length and the effective range are defined by using the effective range expansion (ERE) formula up to the next-leading-order (NLO),
@@ III Two-particle correlation function
The correlation function formulae has been described in many work like [32, 33, 34, 35, 8]. We bring here only the key formulae. The two-particle momentum correlation function for a given potential can be obtained by Koonin-Pratt (KP) formula [35]
@@
The two-particle correlations via KP formula, Eq. (5), is done by employing the ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [36]. CATS tool kit is written in such a way to calculate the correlation function for given an interaction potential and an emission source of any form [37].
The correlation functions utilizing the KP formula, Eq. (5) for three different source sizes, fm and fm are calculated with various potentials and depicted in Fig. 3 a) , b) and c), respectively. The values of source sizes in accord with the typical sizes applied in the exploration of the two-hadron correlation functions [19, 20]. The particular dip shape can be observed at small source size fm, that is conventional in the bound state of system near the threshold [19] (except in the case of QMC potential model by cutoff parameter GeV). Moreover, Fig. 3 (a) shows the results for all cases of two potentials are different at low momentum MeV/c. According to Fig. 1 the HAL potential model, is more attractive than QMC potential model, thus it gives enhancement of .
@@
Where the interaction range is much smaller than source size, it is possible to further investigate the interaction dependence of the correlation function by means of Lednicky-Lyuboshits (LL) formula [38, 35],
with is scattering amplitude and can be obtained by the ERE formula Eq. (4). In addition, and the effective range correction appears in [38, 35]. Since the LL relation is extracted from KP relation by approximating the full wave function by the asymptotic wave function and the effective range correction, by comparing the correlation function via the KP and LL formulae it is possible to decode the detailed shape of potential.
correlation functions is estimated via LL formula (Eq. 6) using the scattering length and the effective range of two models of potentials given in table 1. And, the results are compared with the ones from KP formula in Fig. 4 for three different source sizes and fm. Figure 4 a) and b) show the results of calculations by both potentials HAL (left panel) and QMC (right panel) models for fm source size. They reveal that, the LL approach generates significant different results of the KP formula at low momentum region. In reactions involving light nuclei, the interaction range usually is about fm [13, 12, 39, 19]. Therefore, the LL formula is an incapable approximation where the source size is smaller than the interaction range [35]. Moreover, it can be seen that there is a different behavior about the QMC potential model by cutoff parameter GeV which according to the figure 1 and data in table 1 is the shallowest potential. While the correlation function for another models of potentials have almost same behaviour. So, it is possible to reveal the difference between potentials by precision measurement of correlation function at almost small source size fm. For larger source size fm, the figure 4 e) and f) show that the correlation functions for the LL formula are in acceptable agreement by KP formula ones.
@@ IV Summary and conclusions
In this work, I employed femtoscopy technique to predict momentum correlation functions in the high-energy collisions looking for an additional and alternative source of knowledge relevance to the interaction by using these six different type of potentials. Employing two potentials, correlation functions are calculated using KP formula for three different source sizes, fm and fm. The difference of potentials is appear in the correlation functions by small source size around fm, while for source size fm the correlation functions tended to become same for both potentials. In conclusion, since the correlation functions are sensitive to potentials behavior, we could get important information about effects of particle in dense nuclear medium.
Furthermore, the scattering length and effective range were calculated by solving the Schrödinger equation for each of the potentials separately. Next, correlation functions is examined within the Lednicky-Lyuboshits approach and compared by the results of KP formula. It was seen as expected, the LL formula by small source ( fm) seriously differ from KP formula in the low-momentum region.
@@ Acknowledgement
I am grateful for the authors and maintainers of ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [36], a modified version of which is used for calculations in this work.

### v2

@@ Abstract
Very recently the Wood-Saxon (WS) type interaction in the single-folding potential approach are constructed to simulate the potentials. One set of the potentials are based on the first principle HAL QCD interactions in channel, and in another set, the -meson nucleus potentials were calculated by employing the quark-meson coupling (QMC) model. By utilizing these two set of potentials, the two-particle momentum correlation of in high-energy heavy ion collisions is explored. The numerical results show that the correlation functions at small source size (high density nuclear medium) depends on the employed potential model. Also, the correlation functions are obtained within the Lednicky-Lyuboshits (LL) formalism. For small source size, it is found that, the LL formula returns significantly different values due to the large interaction range of the potential.
@@ I Introduction
Although considerable experimental and theoretical [1, 2, 3, 4, 5, 6, 7] efforts have been done in the last few decades, the possible existence of -mesic bound states with nucleons (N) is still a challenging field of research. The essential basis for the studying of such a system is an exact information about the strong interaction.
@@
Up to now, no evidence of a bound state has been found, experimentally, since the standard invariant mass measurements of its decay products are disputing and the measurement of the scattering parameters are restricted to spin averaged quantities. A list of measurements that have been made for absorption off different nuclear targets is given in [8], and the authors concluded that the landscape of experimental results for the interaction is not convincingly in agreement and new information are required. One method for understanding the hadron-hadron interaction (that is hard to investigate in scattering experiment) is measuring the momentum correlation functions in high-energy collisions [9]. It can provides information on both the effective emission source and the interaction potential. Recent measurement of the correlation function of proton- in heavy-ion collisions by ALICE collaboration [10], together by indicating a bound state using two-particle correlation functions [8, 7, 11] desire the proton bound state hypothesis.
@@
To go one step further, the femtoscopic analysis of hadron-nuclei correlation functions could play a crucial role in understanding the structure and dynamics of the atomic nuclei [12, 13, 14]. Accordingly, the correlation functions are investigated for some hadron-deuteron systems like [15, 13, 16], [17, 16], [18], [19], and [20, 21]. Furthermore, the momentum correlation between [22], [23] and [24] have recently been studied theoretically. From experimental point of view, the momentum correlation between (anti)protons and (anti)deuterons measured in pp collisions at TeV with ALICE is studied for the first time [25] and the first measurement of correlation with GeV Au+Au collisions is done by STAR Collaborations [26]. In addition, there is a research plan at the Thomas Jefferson National Accelerator Facility (JLab), to investigate the binding of (and ) to [4].
@@
A recent HAL QCD simulations of the potential in channel (by the representation that and are the total spin, orbital angular momentum and total angular momentum, respectively) are performed [27]. The simulation is done on the -flavor with nearly physical quark masses MeV on a large lattice volume of at the imaginary-time slices where fm is the lattice spacing. In Ref. [28] the potential was obtained through a single folding of the HAL QCD interaction in the channel with the nuclear matter distribution of . Moreover, a Wood-Saxon (WS) type interaction was built to simulate the potential, given in [4], based on an effective Lagrangian approach where the nuclear density distributions, as well as the in-medium and meson masses, were consistently calculated in the frame work of quark-meson coupling (QMC) model [29]. In Ref. [28], the authors concluded that there were qualitatively a good agreement between binding energies for both types of the interactions.
@@ II Features of potentials
The single folding approach approximates the effective nuclear potential from two body potential, as [30, 31, 32]
where is the nucleon density distribution in -particle at a distance from its center-of-mass. The detailed calculations of how to obtain the is given by Filikhin et al. in Ref. [28]. Here, I just mention the main part. As follows from Ref. [33], Filikhin et al. chose the simple Gaussian matter density distribution model as
@@
In Ref. [28], for phenomenological application and calculation of observables, such as scattering phase shifts and binding energies, the is fitted to a Wood-Saxon form using the function that given by Eq. (3) (motivated by common Dover-Gal model of potential [34]),
where the depth parameter indicates the strength of interaction, where is the mass number of the nucleus and shows the surface diffuseness. The obtained value of the parameters of Eq. (3) correspond to these three rms radius are given in the table 1 and illustrated in Fig. 1. Note that concrete parameterizations are taken straight from Ref. [28].
@@
In Ref. [4], the -meson–nuclear potentials were calculated employing a local density approximation, by evaluating the loop contribution in the self-energy. The nuclear density distributions, as well as the in-medium and meson masses, were consistently calculated in the frame work of quark-meson coupling (QMC) model [29]. Furthermore, they investigated the sensitivity of results to the cutoff parameter in the form factor at the vertex appearing in the -meson self-energy. The resultant potential is given for three values of the cutoff parameter, i.e., and GeV.
Accordingly, Filikhin et al. in Ref. [28], constructed a Wood-Saxon (WS) type potentials to simulate potentials for and GeV. In Fig. 1, for comparison, I just represent the potentials, denoted as by using the parameters given in Ref. [28] directly, for three values of the cutoff parameter and GeV. In table 1, the parameters of potential in Eq. (3) based on HAL QCD (at lattice Euclidean time ) and quark-meson coupling model of potentials, are given.
@@
By employing potentials as input, the Schrödinger equation is solved to extract binding energy and scattering observables from the asymptotic behavior of the wave function. Fig. 2 depicts the normalized phase shifts calculated as functions of the magnitude of the relative momentum ( is the reduced mass of system) with model of HAL QCD for the three values of rms and QMC potentials for the three values of . The phase shift behaviour for the all case represents an attractive interaction even to form a bound state. The scattering length and the effective range are defined by using the effective range expansion (ERE) formula up to the next-leading-order (NLO),
@@ III Two-particle correlation function
The correlation function formulae has been described in many work like [35, 36, 37, 38, 9]. We bring here only the key formulae. Two-particle momentum correlation function for a given potential can be obtained by Koonin-Pratt (KP) formula [38]
@@
The calculations of two-particle correlations via the KP formula, Eq. (5), is done by employing the ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [39]. For given an interaction potential and an emission source CATS tool kit computes the correlation function [40].
The correlation functions by the KP formula, Eq. (5) for three different source sizes, fm and fm are calculated with various potentials and depicted in Fig. 3 a) , b) and c), respectively. The values of source sizes in accord with the typical sizes applied in the exploration of two-hadron correlation functions [22, 23]. The particular dip shape can be observed at small source size fm, that is conventional in the bound state of system near the threshold [22] (except in the case of QMC potential model by cutoff parameter GeV). Moreover, Fig. 3 (a) shows the results for all cases of two potentials are different at low momentum MeV/c. According to Fig. 1 the HAL potential model, is more attractive than QMC potential model, thus it gives enhancement of .
@@
In the two-body systems if the interaction range is much smaller than source size, it is possible to further investigate the interaction dependence of the correlation function by means of Lednicky-Lyuboshits (LL) formula [41, 38],
is scattering amplitude and it is evaluated from the ERE formula Eq. (4). In addition, and [41, 38] are mathematical functions. Since the LL relation is extracted from the KP relation by approximating the full wave function by the asymptotic wave function and the effective range correction, by comparing the correlation function via the KP and the LL formulae it is possible to decode the detailed shape of potential.
correlation functions is estimated via the LL formula (Eq. 6) using the scattering length and the effective range of two models of potentials given in table 1. And, the results are compared with the ones from the KP formula in Fig. 4 for three different source sizes and fm. Figure 4 a) and b) show the results of calculations by both potentials HAL (left panel) and QMC (right panel) models for fm source size. They reveal that, the LL approach generates significant different results of the KP formula at low momentum region. In reactions involving light nuclei, the interaction range usually is fm [16, 15, 42, 22]. Therefore, the LL formula is an incapable approximation where the source size is smaller than the interaction range [38]. Moreover, it can be seen that there is a different behavior about the QMC potential model by cutoff parameter GeV which according to the figure 1 and data in table 1 is the shallowest potential. While the correlation function for another models of potentials have almost same behaviour. So, it is possible to reveal the difference between potentials by precision measurement of correlation function at almost small source size fm. For larger source size fm, the figure 4 e) and f) show that the correlation functions for the LL formula are in acceptable agreement by the KP formula ones.
@@ IV Summary and conclusions
In this work, I employed femtoscopy technique to predict momentum correlation functions in the high-energy collisions looking for an additional and alternative source of knowledge relevance to the interaction by using these six different type of potentials. Employing two potentials, correlation functions are calculated using the KP formula for three different source sizes, fm and fm. The difference of potentials is appear in the correlation functions by small source size around fm, while for source size fm the correlation functions tended to become same for both potentials. In conclusion, since the correlation functions are sensitive to potentials behavior, we could get important information about effects of particle in dense nuclear medium.
Furthermore, the scattering length and effective range were calculated by solving the Schrödinger equation for each of the potentials separately. Next, correlation functions is examined within the Lednicky-Lyuboshits approach and compared by the results of the KP formula. It was seen as expected, the LL formula by small source ( fm) seriously differ from the KP formula in the low-momentum region.
@@ Acknowledgement
I am grateful for the authors and maintainers of ”Correlation Analysis tool using the Schrödinger Equation” (CATS) [39], a modified version of which is used for calculations in this exploratory study. Discussions during the long-term workshop, HHIQCD2024 at Yukawa Institute for Theoretical Physics (YITP-T-24-02), were useful as I finished this work.